# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

### v2

@@ 4.2. Case 2: $n\in\{2,3\}\quad\&\quad q\in\left[\frac{n+3}{n-1},~\frac{n+2}{(n-2)_{+}}\right)$
$V$ $\geq V_{0}-\widetilde{M}\bar{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
$W$ $\geq W_{0}-\widetilde{M}\bar{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
@@
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},~+\infty\right),$
@@ 5. Geometric applications
Now we want to make use of some conclusions obtain in the previous sections to study the topological properties of a manifold which enjoy a Sobolev inequality. As mentioned in the previous, many manifolds, for instances Riemannian manifold with ${\rm Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, minimal hypersurfaces [43] and Ricci shrinker [39], support Sobolev inequalities. In particular, Brendle [7] obtained the sharp Sobolev inequality on complete noncompact Riemannian manifolds with nonnegative Ricci curvature using the so-called ABP method, addressing an open question posed by Cordero-Erausquin, Nazaret, and Villani [21] for such manifold. Balogh and Kristály [3] provided an alternative proof of Brendle’s rigidity result and confirmed that the Sobolev constant is sharp. Concretely, their results can be stated as follows:
Let $\left(M^{n},g\right)$ be a noncompact, complete $n$-dimensional Riemannian manifold with ${\rm Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, where $\rm Ric$} denote the Ricci curvature of $(M,g)$. Then for all $v\in C_{0}^{\infty}\left(M^{n}\right)$, there holds
@@
$M\setminus\bar{U}=\cup_{i=1}^{k}E_{i}$
with each $E_{i}$ being a noncompact connected component of $M\setminus\bar{U}$. Usually, we also call $E_{1},\cdots,E_{k}$ as the ends corresponding to $U$.
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold with $\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and ${\rm Ric}(M)\geq 0$ outside some compact subset. Then $M$ has only finitely many ends $E_{1},E_{2},\cdots,E_{k}$ and $\dim H^{\infty}(M)=k$.
In fact, this theorem and the Cheng-Yau’s gradient estimate of positive harmonic functions (see [16]) imply that a complete noncompact Riemannian manifold, which satisfies $n=\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and ${\rm Ric}(M)\geq 0$, has only an end.
@@
We let $\Omega_{i}$ ($i=1,2,\cdots$) be an exhaustion of $(M,g)$, i.e., each $\Omega_{i}$ is an open domain contained in $M$ and $\bar{\Omega}_{i}$ is compact, $\bar{\Omega}_{i}\subset\Omega_{i+1}$ for every $i\geq 1$, and $\cup_{i=1}^{\infty}\Omega_{i}=M$. For $i\geq i_{0}$ and $i_{0}$ sufficiently large, let
@@
be the disjoint union of connected components. Denote the $k$ ends of $(M,g)$ by $E_{1}^{(i_{0})},\cdots,E_{k-1}^{(i_{0})}$ and $E_{k}^{(i_{0})}$, and for the sake of convenience, by $E_{1},\cdots,E_{k}$ without confusions.
@@
As done in [26], we let $\phi_{j}(x)$ be a real smooth function on $(M,g)$ such that
$\phi_{j}(x)\equiv 1\quad\mbox{on}\,\,E_{j}^{(i_{0})}\quad\mbox{and}\quad\phi_{j}(x)\equiv 0\quad\mbox{in}\,\,M\setminus(\Omega_{i_{0}}\cup E_{j}^{(i_{0})});$
@@
$\rho_{j,i}=|u^{i}_{j}-\phi_{j}|\quad\mbox{for}\,\,i\geq i_{0}.$
It is easy to see that $|\Delta\phi_{j}|\in L^{p}$ for any $p>0$ and in the sense of distribution (see [26])
$\Delta\rho_{j,i}\geq-|\Delta\phi_{j}|.$
@@
$\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}\Delta\rho_{j,i}\geq-\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\phi_{j}|.$
@@
$\frac{4(\alpha-1)}{\alpha^{2}}\int_{\Omega_{i}}\left|\nabla\rho_{j,i}^{\frac{\alpha}{2}}\right|^{2}\leq\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\phi_{j}|.$
@@
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq\frac{\alpha^{2}}{4(\alpha-1)}\mathbb{S}_{\chi}(M)^{-1}\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\phi_{j}|.$
@@
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq C\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{\alpha-1}{\alpha\chi}}\left(\int_{\Omega_{i}}|\Delta\phi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$
@@
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}\leq C\left(\int_{\Omega_{i}}|\Delta\phi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$ (5.1)
@@
$\left(\int_{E_{j}}(1-u_{j})^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\quad\mbox{and}\quad\quad\left(\int_{E_{l}}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\mbox{for}\,\,l\neq j,$ (5.2)
@@
$\int_{M}|\Delta\phi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}<\infty.$
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
Let $(M^{n},g)$ be a complete noncompact Riemannian $n$-dimensional manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$ where $1<\chi\leq\frac{n}{n-2}$. If ${\rm{vol}}(B_{r}(x_{0}))=O(r^{\frac{2\chi}{\chi-1}})$ for geodesic ball $B_{r}(x_{0})\subset M$ where $x_{0}\in M$ is a fixed point, then $M$ has only finitely many ends. In particular, in the case $\chi=\frac{n}{n-2}$ and ${\rm{vol}}(B_{r})=O(r^{n})$ for any geodesic ball $B_{r}\subset M$, $M$ has only finitely many ends.
@@
where $q=\#\{x_{j}\}$. In fact, if there exists some $\bar{x}\in\partial B_{r}(x_{0})\setminus\cup_{i=1}^{q}B_{\frac{r}{2}}(x_{i})$, then $B_{\frac{r}{4}}(\bar{x})$ is disjoint with $B_{\frac{r}{4}}(x_{i})$ for $i=1,\cdots,q$. This contradicts the fact that $\{x_{i}\}$ is a maximal family.
@@
where $C(n)/C\left(\chi,\mathbb{S}_{\chi}(M)\right)]$ is the integer part of $C(n)/C\left(\chi,\mathbb{S}_{\chi}(M)\right)$. This means that the counting number of the above maximal family denoted by $\{x_{j}\}$ is bounded by a constant depending only on $n$, $\chi$ and $\mathbb{S}_{\chi}(M)$.
@@
be the disjoint union of connected components.
Now, we pick $x_{j}^{\prime}\in\partial B_{4r_{0}}(x_{0})\cap E^{(r_{0})}_{j}$ (where $j=1,\cdots,s$), then, it is easy to see that $B_{r_{0}}(x_{j}^{\prime})$ ($j=1,\cdots,s$) are disjoint with each other. Indeed, if $B_{r_{0}}(x_{i}^{\prime})\cap B_{r_{0}}(x_{j}^{\prime})$ is not a empty set where $i\neq j$, then we know $d(x_{i}^{\prime},x_{j}^{\prime})<2r$. On the other hand, there a minimal geodesic connecting $x_{i}^{\prime}$ and $x_{j}^{\prime}$, denoted by $\sigma(s)$, which must pass through $\partial B_{r_{0}}(x_{0})$, since $x_{i}^{\prime}$ and $x_{j}^{\prime}$ belong to different connected components. Hence $d(x_{i}^{\prime},x_{j}^{\prime})>3r$, this is contradiction.
But we take $\Lambda$ geodesic balls, denoted by $B_{2r_{0}}(x_{j_{i}}^{\prime})$ where $\{j_{i}:i=1,\cdots,\Lambda\}\subset\{1,2,\cdots,s\}$, it is easy to see the union $\cup_{i=1}^{\Lambda}B_{2r_{0}}(x_{j_{i}}^{\prime})$ can not cover $\partial B_{4r_{0}}(x_{0})$. This is a contradiction. ∎
Acknowledgments: Y. Wang is supported by National Natural Science Foundation of China (Grant No.12431003); G. Wei is supported by National Natural Science Foundation of China (Grants No.12101619 and 12141106).

### v3

@@ 4.A.2. Integral estimate
$V$ $\geq V_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
$W$ $\geq W_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
@@
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},\penalty 10000\ +\infty\right),$
@@ 5. Geometric applications
Now we want to make use of some conclusions obtain in the previous sections to study the topological properties of a manifold which enjoy a Sobolev inequality. As mentioned in the previous, many manifolds, for instances Riemannian manifold with $\mathrm{Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, minimal hypersurfaces [49] and Ricci shrinker [45, 66], support Sobolev inequalities. In particular, Brendle [8] obtained the sharp Sobolev inequality on complete noncompact Riemannian manifolds with nonnegative Ricci curvature using the so-called ABP method, addressing an open question posed by Cordero-Erausquin, Nazaret, and Villani [24] for such manifold. Balogh and Kristály [4] provided an alternative proof of Brendle’s rigidity result and confirmed that the Sobolev constant is sharp. Concretely, their results can be stated as follows:
Let $\left(M^{n},g\right)$ be a noncompact, complete $n$-dimensional Riemannian manifold with $\mathrm{Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, where $\mathrm{Ric}$ denote the Ricci curvature of $(M,g)$. Then for all $v\in C_{0}^{\infty}\left(M^{n}\right)$, there holds
@@
$M\setminus\overline{U}=\cup_{i=1}^{k}E_{i}$
with each $E_{i}$ being a noncompact connected component of $M\setminus\overline{U}$. Usually, we also call $E_{1},\cdots,E_{k}$ as the ends corresponding to $U$.
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold with $\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and $\mathrm{Ric}(M)\geq 0$ outside some compact subset. Then $M$ has only finitely many ends $E_{1},E_{2},\cdots,E_{k}$ and $\dim H^{\infty}(M)=k$.
In fact, this theorem and the Cheng-Yau’s gradient estimate of positive harmonic functions (see [17]) imply that a complete noncompact Riemannian manifold, which satisfies $n=\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and $\mathrm{Ric}(M)\geq 0$, has only an end.
@@
We let $\Omega_{i}$ ($i=1,2,\cdots$) be an exhaustion of $(M,g)$, i.e., each $\Omega_{i}$ is an open domain contained in $M$ and $\overline{\Omega}_{i}$ is compact, $\overline{\Omega}_{i}\subset\Omega_{i+1}$ for every $i\geq 1$, and $\cup_{i=1}^{\infty}\Omega_{i}=M$. For $i\geq i_{0}$ and $i_{0}$ sufficiently large, let
@@
be the disjoint union of connected components. Pick $k$ ends of $\{E^{(i)}_{j}:1\leq j\leq s\}$ and denote the $k$ ends of $(M,g)$ by $E_{1}^{(i_{0})},\cdots,E_{k-1}^{(i_{0})}$ and $E_{k}^{(i_{0})}$, and for the sake of convenience, by $E_{1},\cdots,E_{k}$ without confusions.
@@
As done in [30], we let $\varphi_{j}(x)$ be a real smooth function on $(M,g)$ such that
$\varphi_{j}(x)\equiv 1\quad\mbox{on}\,\,E_{j}^{(i_{0})}\quad\mbox{and}\quad\varphi_{j}(x)\equiv 0\quad\mbox{in}\,\,M\setminus(\Omega_{i_{0}}\cup E_{j}^{(i_{0})});$
@@
$\rho_{j,i}=|u^{i}_{j}-\varphi_{j}|\quad\mbox{for}\,\,i\geq i_{0}.$
It is easy to see that $|\Delta\varphi_{j}|\in L^{p}$ for any $p>0$ and in the sense of distribution (see [30])
$\Delta\rho_{j,i}\geq-|\Delta\varphi_{j}|.$
@@
$\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}\Delta\rho_{j,i}\geq-\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\varphi_{j}|.$
@@
$\frac{4(\alpha-1)}{\alpha^{2}}\int_{\Omega_{i}}\left|\nabla\rho_{j,i}^{\frac{\alpha}{2}}\right|^{2}\leq\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\varphi_{j}|.$
@@
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq\frac{\alpha^{2}}{4(\alpha-1)}\mathbb{S}_{\chi}(M)^{-1}\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\varphi_{j}|.$
@@
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq C\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{\alpha-1}{\alpha\chi}}\left(\int_{\Omega_{i}}|\Delta\varphi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$
@@
(5.1) $\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}\leq C\left(\int_{\Omega_{i}}|\Delta\varphi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$
@@
(5.2) $\left(\int_{E_{j}}(1-u_{j})^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\quad\mbox{and}\quad\quad\left(\int_{E_{l}}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\mbox{for}\,\,l\neq j,$
@@
$\int_{M}|\Delta\varphi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}<\infty.$
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
Let $(M^{n},g)$ be a complete noncompact Riemannian $n$-dimensional manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$ where $1<\chi\leq\frac{n}{n-2}$. If ${\rm{Vol}}(B_{r}(x_{0}))=O\left(r^{\frac{2\chi}{\chi-1}}\right)$ for some geodesic ball $B_{r}(x_{0})\subset M$, then $M$ has only finitely many ends.
In particular, suppose $(M^{n},g)$ supports a $\chi$-type Sobolev inequality for some $\chi\in(1,n/(n-2))$, and assume in addition that $\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}$ is bounded. Then $M$ has only finitely many ends.
@@
where $q=\sharp\{x_{j}\}$. In fact, if there exists some $\overline{x}\in\partial B_{r}(x_{0})\setminus\cup_{i=1}^{q}B_{\frac{r}{2}}(x_{i})$, then $B_{\frac{r}{4}}(\overline{x})$ is disjoint with $B_{\frac{r}{4}}(x_{i})$ for $i=1,\cdots,q$. This contradicts the fact that $\{x_{i}\}$ is a maximal family.
@@
where $C(n)/C\left(\chi,\mathbb{S}_{\chi}(M)\right)]$ is the integer part of $C(n)/C\left(\chi,\mathbb{S}_{\chi}(M)\right)$. This means that the counting number of the above maximal family denoted by $\sharp\{x_{j}\}$ is bounded by a constant depending only on $n$, $\chi$ and $\mathbb{S}_{\chi}(M)$.
@@
is the disjoint union of connected components.
Now choose points $x_{j}^{\prime}\in\partial B_{4r_{0}}(x_{0})\cap E^{(r_{0})}_{j}$ for $j=1,\dots,s$. It is then straightforward to see that the balls $B_{r_{0}}(x_{j}^{\prime})$ $(j=1,\dots,s)$ are pairwise disjoint. Indeed, if $B_{r_{0}}(x_{i}^{\prime})\cap B_{r_{0}}(x_{j}^{\prime})\neq\varnothing$ for some $i\neq j$, then $d(x_{i}^{\prime},x_{j}^{\prime})<2r_{0}$. On the other hand, since $x_{i}^{\prime}$ and $x_{j}^{\prime}$ lie in different connected components, any minimizing geodesic joining them must intersect $\partial B_{r_{0}}(x_{0})$. Hence $d(x_{i}^{\prime},x_{j}^{\prime})\geq 6r_{0},$ a contradiction. Therefore the family $\{B_{r_{0}}(x_{j}^{\prime})\}_{j=1}^{s}$ is disjoint, which contradicts the fact that the cardinality of maximal disjoint family is at most $\Lambda$. We complete the proof. ∎
Acknowledgments: We would like to thank Professor Jun Yang for insightful discussions. Y. Wang is supported by National Natural Science Foundation of China (Grant No.12431003); G. Wei is supported by National Natural Science Foundation of China (Grants No.12101619 and 12141106).